11. Two years ago the Krusty Krab Restaurant purchased a grill for $50,000. The owner, Eugene Krabs, has learned that a new grill is available that will cook Krabby Patties twice as fast as the existing grill. This new grill can be purchased for $80,000 and would be depreciated straight line over 8 years, after which it would have no salvage value. Eugene Krab expects that the new grill will produce EBITDA of $50,000 per year for the next eight years while the existing grill produces EBITDA of only $35,000 per year. The current grill is being depreciated straight line over its useful life of 10 years after which it will have no salvage value. All other operating expenses are identical for both grills. The existing grill can be sold to another restaurant now for $30,000. The Krusty Krab's tax rate is 35%. If the Krusty Krab's opportunity cost of capital is 12%, then the NPV for upgrading to the new grill is closest to:

Answer is 10,630 - just looking for the steps to solve. Thanks!
Answer #1
1] INITIAL INVESTMENT:
Cost of the new grill $          80,000
-After tax sale value of old grill:
Current sale value of the old grill $           30,000
Book value of the old grill = 50000-(50000/10)*2 = $           40,000
Loss on sale of old grill = 40000-30000 = $           10,000
Tax shield at 35% = 10000*35% = $             3,500
After tax sale value of old grill = 30000+3500 = $          33,500
Initial investment $          46,500
2] INCREMENTAL OCF:
Incremental EBITDA = 50000-35000 = $           15,000
-Incremental depreciation = 80000/8-50000/10 = $             5,000
=Incremental EBIT $           10,000
-Tax at 35% $             3,500
=Incremental NOPAT $             6,500
+Depreciation $             5,000
=Incremental OCF $           11,500
3] CALCULATION OF NPV:
PV of incremental OCF = 11500*PVIFA(12,5) = 11500*4.96764 $           57,128
Less: Initial investment $           46,500
NPV $           10,628
Can be rounded off to $10,630.